OpenStudy (anonymous):

A spherical snowball is melting in such a way that its diameter is decreasing at rate of 0.3 cm/min. At what rate is the volume of the snowball decreasing when the diameter is 16 cm.

OpenStudy (anonymous):

120.6371579 possibly

OpenStudy (anonymous):

Bring back memories huh, mnad1:)

OpenStudy (anonymous):

haha yeah, deja vu

OpenStudy (bahrom7893):

V = (4/3)pir^3 dv/dt = 4pir^2(dr/dt)

OpenStudy (bahrom7893):

d = 16, so r = 8 dv/dt = 4pi * 8^2 * 0.3

OpenStudy (anonymous):

differntiate the formular for volume of a sphere, then use 8 for the radius and 0.15 for dd/dt(meaing the rate of change of the diameter)

OpenStudy (bahrom7893):

dv/dt = 76.8pi

OpenStudy (bahrom7893):

dv/dt = 241.152...

OpenStudy (bahrom7893):

someone confirm/reject my answer.. i am coming back from a math exam.. my brain is dead tired.. and im doin this off the top of my head..

OpenStudy (bahrom7893):

oh wait I AM SOO WRONG!

OpenStudy (bahrom7893):

A spherical snowball is melting in such a way that its diameter is decreasing at rate of 0.3 cm/min. At what rate is the volume of the snowball decreasing when the diameter is 16 cm.

OpenStudy (anonymous):

you're not. your ans/2 =correct.

OpenStudy (anonymous):

the answer 120.637 is correct

OpenStudy (anonymous):

okay... V = 4/3*pi*r^3 = 4/3*pi*(d/2)^3 = 4/3*pi*d^3/8 = 1/6*pi*d^3 dV/dt = 1/2*pi*d^2*dd/dt dV/dt = -0.3, when d = 16: -0.3 = 0.5*pi*256*dd/dt so dd/dt = -0.3/(128*pi) so decreasing at 0.3/(128*pi)

OpenStudy (anonymous):

dv/dt=(4)(pi)(8^2)(0.15)=120.637

OpenStudy (anonymous):

oh...i got the diameter decrease rate... lol dv/dt = 0.5*pi*256*-0.3 = -120.64 so decreasing at 120.64

OpenStudy (bahrom7893):

Okay I got -38.4pi, which is 120.64 I confirm..

OpenStudy (bahrom7893):

TOGETHER WE SHALL DEFEAT RELATED RATES!

OpenStudy (bahrom7893):

sorry meant which is -120.64 lol

OpenStudy (bahrom7893):

All of yall.. good answers everyone xD!

OpenStudy (anonymous):

good effort